How many three-digit numbers can you form from the digits 1, 2, 3, 4, 5, 6 if:
a.) the digits can be repeated
b.) The digits can not be repeated

Answer by stanbon(75887)

How many three-digit numbers can you form from the digits 1, 2, 3, 4, 5, 6 if:
a.) the digits can be repeated
6^3 = 216
--------------
b.) The digits can not be repeated
# of ways to pick three distanct numbers: 6C3 = [6*5*4]/[1*2*3] = 20
# of ways to arrange a particular group of three digits: = 3!=6
Total number of ways = 20*6 = 120 ways
==========
Cheers,
Stan H.